OpenStudy (anonymous):

how is the equation written? Help! I need to understand the following:Write an equation of the line parallel to the x-axis and passing through (3, 5)

OpenStudy (anonymous):

A line that is parallel to x-axis has a slope of 0. The line equation form is \[y-y1=m(x-x1)\]. So, you have m=0, and plugs in the value \[y-5=0(x-3)\]; final answer is y=5. Check out more tips on http://www.oojih.com/show/algebra/linearfunction/

OpenStudy (amistre64):

i spose first you need to know what an xaxis is; then what it means to be parallel;....